A 2kg block at rest is acted upon by a constant force of 20N for 4 seconds. Find the final velocity of the block. Show complete solutions.




1
Expert's answer
2022-01-12T08:29:49-0500

Given

Mass, m=2 kg

Force, F=20 N

Time, t=4 s


Find

The final velocity of the block v


Solution

By Newton's second law in terms of momentum change, we remember that


"I=\Delta p=m\Delta v=m(v-v_i),\\\space\\
\Delta p=Ft,\\\space\\
m(v-v_i)=Ft,\\\space\\
2(v-0)=20·4,\\\space\\
v=40\text{ m/s}."